OpenStudy (anonymous):

Find the distance between (-2, 8) and (11, 2). Round to the nearest hundredth.

OpenStudy (lgbasallote):

\[\text{distance} = \sqrt{(x_2 - x_1)^2 + (y_2 - y_1)^2}\] \[\text{distance} = \sqrt{(-2 - 11)^2 + ( 8 - 2)^2}\] so what's your answer here?

OpenStudy (lgbasallote):

@elliothoo did you get that?